OpenStudy (anonymous):

Find the maximum or minimum of the following quadratic function: y = -10x^2 + x - 10. A. -40 B. -10 C. 10 D. -399/40

OpenStudy (mathmale):

Just by looking at the first term of this function, you can tell whether the graph opens UP or DOWN. Which is it, and why?

OpenStudy (anonymous):

It's in the negatives

OpenStudy (mathmale):

I don't understand. Explain again, please. Next, Would you please use x=-b/(2a) to find the x-coordinate of the vertex of your parabola? x = ?

OpenStudy (anonymous):

i got the the answer its -10

OpenStudy (anonymous):

i graphed it.

OpenStudy (mathmale):

that's a great way to learn what's happening here! Congrats. Note that if you had found x=-b/(2a), you'd have gotten x=-(1)/(2*[-10]) = 1/20, and using this x value would give you the corresponding y value.

OpenStudy (mathmale):

Note that you still have to say whether you've found a max or a min.